The length and width of a rectangular field are in the ratio 4:3. If the area is 432 square units, find its perimeter.

What I have done so far was multiply 4 by 3 and divided the product by 432.

Answer by rfer(16322)

4x*3x=432
12x^2=432
x^2=432/12
x^2=36
x=sqrt 36
x=6
4x=24
3x=18
-------------
P=2(24)+2(18)
P=48+36
P=84 units